# Applied-field magnetic structure and spectroscopy shifts of the effective spin-, -like magnet Li2CoCl4

### v1

@@ Abstract
Insulators containing chains of magnetic transition metal cations provide platforms for probing spin- dynamics and quantum critical behavior. Li2CoCl4 contains edge-sharing CoCl6 octahedra that form chains along the crystallographic axis and orders antiferromagnetically at zero field, but questions remain about its applied-field magnetic structure and the Co2+ spin state. Here, we show with neutron diffraction on a polycrystalline sample how the anti-aligned chains of cobalt moments undergo a spin-flop transition to a field-aligned ferromagnetic state above 1.6 T. Further, using magnetic resonance absorption measurements and paramagnetic spin models, we reveal the strongly anisotropic nature of the Co2+ ion’s -like magnetic behavior ( and ) and its ground state. We, therefore, supply the magnetic structures and anisotropic description needed to explore the dynamics of the field-driven magnetic phases, laying the foundation for further experimental and theoretical studies.
@@ I Introduction
At zero field, Li2CoCl4 orders antiferromagnetically with magnetic space group symmetry. The Co2+ magnetic moments align within edge-sharing CoCl6 chains and anti-align between nearest neighbor chains. The magnetic susceptibility and heat capacity shift with increasing field, leading to three proposed magnetic phase regions separated by transitions near 1.6 T and 3.5 T at 2 K [18]. The magnetic structures of the applied-field regions are unknown. The low-field region is antiferromagnetic, likely similar to the zero-field magnetic ordering. The intermediate-field region may be ferrimagnetic, with chain rotations reminiscent of the structurally similar compound CoCl2H2O [23, 24], though the data were also consistent with a spin-flop reorientation. Finally, in the high-field region, not sharply delineated from the intermediate-field one, the Li2CoCl4 moment chains are likely field-aligned, with no well-defined Curie temperature in the magnetic susceptibility.
Here, we use polycrystalline Li2CoCl4 to identify the applied-field regions’ magnetic structures with neutron diffraction, particularly that of the previously ambiguous intermediate-field region. Additionally, we extract several parameters from applied-field infrared spectroscopy data, such as the anisotropic -factors, to discuss the role of spin-orbit coupling and anisotropic interactions on the magnetic properties of Li2CoCl4.
@@ II Materials and methods
We collected neutron diffraction data on the HB-2A beamline at the High Flux Isotope Reactor at Oak Ridge National Laboratory. To avoid grain reorientation under applied field, we pressed pellets of Li2CoCl4 for the diffraction experiment. Under argon, we pressed 1.5 g of powder into around 30 1/8” diameter pellets. We then sealed the pellets under vacuum, transported them to ORNL, and loaded them under helium into a 6 mm diameter vanadium canister. The top of the canister was packed with aluminum foil before being hermetically sealed.
@@
We performed magneto-infrared measurements at the infrared (IR) spectroscopy facility at the National High Field Magnetic Laboratory with a 17.5 T vertical-bore superconducting magnet coupled to a Bruker Vertex 80v FTIR spectrometer. In an argon glovebox, a 3–5 mg polycrystalline sample was bonded with n-eicosane and loaded between two polypropylene layers to protect it from oxygen and moisture. We placed the sample in a Voigt geometry so that the incoming IR light was perpendicular to the magnetic field. A composite Si bolometer (IR Labs) collected the transmitted IR light with a spectral range of 10–1100 cm-1 (0.3–33 THz) and a resolution of 0.3 cm-1 (9 GHz). Additionally, we placed a low-pass THz filter (QMC Instruments) with a 120 cm-1 cutoff frequency in front of the sample to increase sensitivity in the far-IR (FIR) range and to minimize radiative heating. IR data were collected near 5 K.
To isolate the magnetic resonance absorption, the spectra were divided by the average spectrum for all magnetic fields. Such normalized spectra are only sensitive to intensity shifts due to magnetic field and eliminate contributions from nonmagnetic vibrational absorption and instrument features. Our data analysis utilized the EPR analysis package EasySpin [28].
III Results
@@ III.1 Zero- and low-field magnetic structures
We first confirmed that the zero-field magnetic structure matches that determined previously. The data refines well to the antiferromagnetic structure with (BNS no. 55.363) magnetic space group symmetry (see supplemental material [29]). For the magnetic space group , only reflections where are allowed, whereas the nuclear space group only allows reflections where , making it simple to identify the magnetic ordering reflections. The magnetic structure is commensurate, with the propagation vector k = (1,0,0). The magnetic moments are aligned within the Co2+ chains and anti-aligned between nearest neighboring chains along the nuclear cell’s -plane diagonal. The refined moment is 2.10(3) . The 20 K, zero-field data in the paramagnetic region is included in the supplemental material for comparison [29].
@@
Prior magnetic susceptibility and heat capacity data indicated that the low-field magnetic structure ( 1.6 T) matches the zero-field structure. At 0.8 T, we observe the same magnetic reflections as in the zero-field case, associated with a loss of -centering, but with slightly decreased intensity. Refinements give an unconstrained magnetic moment of 2.02(3) (Fig. 1), indicating some magnetic intensity may become paramagnetic background, though restricting the moment to 2.10 gives a nearly identical fit. With the field increased to 1.5 T, just below the 1.6 T transition to the intermediate-field phase, we can again refine the data well to the structure with a constrained 2.10 moment or a refined moment of 1.96(3) . See the supplemental material [29] for the 1.5 T refinement.
@@ III.2 Field-induced moment rotation
Between the 1.5 T and 2.5 T scans, the antiferromagnetic structure reflections, e.g. (010), (100), and (120), lose significant intensity while the reflections allowed by the -centered nuclear cell, e.g. (110) and (020), gain intensity. The intensity shifts indicate a spin reorientation to a k = 0 magnetic structure. Unlike for CoCl2H2O [23, 24], we do not observe new peaks associated with a ferrimagnetic phase. Instead, a spin-flop transition begins to align the chains of Co2+ magnetic moments.
Because the measured sample is polycrystalline, the magnetic field is not along a specific crystallographic direction and the spin-flop transition is spread out, leading to a slow evolution within the sample and a mixture of magnetic phases. Therefore, the best refinement models above 1.6 T involve both the antiferromagnetic magnetic phase and a ferromagnetic k = 0 structure with moments along the crystallographic axis. The moment is constrained to be equal in the two magnetic phases when both are refined. The full range refinements for the 2.5, 3.5, 4.5, and 5.5 T data are shown in Fig. S5 [29]. Fig. 2 shows both magnetic unit cells along with the nuclear cell. The ferromagnetic structure has symmetry (BNS no. 65.485). This maximal magnetic space group only allows a Co2+ moment along .
Several factors oppose modeling the magnetic structure with the other maximal groups (moments along only) and (moments along only). First, the increasing intensity on the (020) reflection violates the systematic absence of (0k0) reflections. Second, at higher fields, the significant increase in intensity on the (110) reflection (Fig. 3), though allowed for both the and symmetries, suggests a preferred moment alignment along . Only the moment component within the (110) plane, and thus perpendicular to the corresponding scattering vector, will contribute to the magnetic intensity. For moments along , this component equals the magnitude of the total moment for the symmetry structure but is less than the magnitude of the total moment for the symmetry structure. Consequently, fitting with does not account for the full (110) reflection intensity increase (Fig. 3). Third, a model with moments along fits nicely with the previously established zero-field structure and the proposed spin-flop transition. At 5.5 T, a majority of the moments are likely aligned along .
Additionally, commensurate magnetic structures that allow for canting do not improve the magnetic peak fits. For example, the 2 space group symmetry (BNS no. 12.62) is the highest symmetry k = 0 magnetic space group possible for Li2CoCl4 that allows for moment components along and . Refinements to the high-field data using 2 ( = 1.5(2) , = 1.7(1) ) or symmetry gave a similar fit [29]. Therefore, we tentatively retain the simpler model of Co2+ moments flipping along with increasing field.
The refinement results for the base temperature scans with the and magnetic space group structures are recorded in Table 1. Refinements to the 2.5 and 5.5 T data, corresponding to the previously described intermediate-field and high-field regions, are in Fig. 4, which highlights the changing magnetic intensity contributions of the antiferromagnetic and ferromagnetic phases. At 2.5 and 3.5 T, the refined magnetic moment dips significantly. The shift corresponds with the antiferromagnetic phase (010) reflection’s drop with increasing field (Fig. 5), with the shift to a steeper slope in the magnetization curve [18], and with emerging phase magnetic intensity. If we constrain the moment for every field to the zero-field value of 2.10 , the 2.5 T refinement shows the largest change [29] while the other results are qualitatively identical.
@@
While tracking the (010) reflection intensity with increasing field, we additionally captured changes in several lower intensity reflections with the other diffractometer detectors: the antiferromagnetic (100) reflection and the nuclear/ferromagnetic (131) reflection. Both show the same general trend of increasing ferromagnetic intensity and decreasing antiferromagnetic intensity above 1.5–1.6 T [29]. Additionally, we monitored the ferromagnetic (110) reflection’s changes with increasing temperature in the high-field region at 5.6 T (Fig. 6). Up to 20 K, we do not observe a sharp drop or leveling off of the intensity but rather a nearly linear decrease in reflection intensity. Similarly, the magnetic susceptibility at high field, even up to 7 T [29], does not show the typical sharp rise of a paramagnetic to ferromagnetic transition, suggesting that the high-field magnetic ordering is driven by field overcoming the material’s interchain antiferromagnetic exchange interactions instead of by dominant interchain ferromagnetic exchange interactions.
III.3 IR spectroscopy in an applied magnetic field
The normalized FIR magnetic resonance absorption contains a broad continuum and a narrow peak (Fig. 7a). The broad continuum shows the strongly anisotropic response of the Li2CoCl4 sample, which contains crystallites randomly oriented relative to the magnetic field. At high fields, where Zeeman effects dominate, we extracted anisotropic -factors of = 2.77 and = 5.23, where and are relative to the compressed CoCl6 axis (crystallographic axis at base temperature). A narrow absorption peak is also observed in the middle of the continuum that is consistent with a -factor of 4.16. The corresponding resonance lines confine the magnetic absorption spectrum for a free cobalt ion, but the experimental data diverge from the lines below 10 T where exchange interactions produce collective spin excitations.
Approaching zero field, the magnetic anisotropy decreases, and the continuum collapses to the same shape for each crystallite. The magnetic field where the absorption transitions between continuum and gapped behavior, notably, matches that of the transition between regions dominated by field-aligned moments and spin-flop behavior. At zero field, we attribute the strong peak at 11.6 cm-1 and the weaker satellite at 17.2 cm-1 (Fig. 7c) to spin gaps in the collective excitations of Li2CoCl4. Because we are unaware of structural analogues to Li2CoCl4 with which we could create a magnetically dilute system, we were unable to probe the 10 cm-1 region with electron paramagnetic resonance measurements.
IV Discussion
IV.1 Magnetic structure evolution
The zero-field magnetic ordering determined by neutron diffraction data agrees with data collected previously [18]. Below and at zero field, Co2+ moments with magnitude 2.10(3) are aligned within chains while antiferromagnetic interactions between chains lead to a cusp in the susceptibility and a negative Curie-Weiss temperature. Though the moment in this study is slightly lower than the previously refined value 2.19(4) , the refinement standard errors are likely underestimated. Since we were able to produce a closer fit to peak intensities in this work, especially nuclear intensity, we will use 2.10 here.
With increasing field, the field-aligned ferromagnetic phase emerges. The notable appearance of the ferromagnetic phase near 1.6 T in neutron data matches the slope increase of the magnetization, the susceptibility’s leveling off below the critical temperature, and the emergence of a hump in the magnetic heat capacity alongside a sharper peak. The suppression of the heat capacity antiferromagnetic transition peak along with the emergence of the heat capacity hump follows from the reduction of antiferromagnetic regions and the progressive aligning of chains in newly formed ferromagnetic regions in the polycrystalline samples. Since the 2.5 T, 1.5 K neutron diffraction data captures the region of greatest competition between the antiferromagnetic and ferromagnetic interactions, which reduces the refined magnetic moment, it is the only pattern where constraining the moment to the zero-field value of 2.10 significantly alters the refinement. At higher fields where the ferromagnetic phase dominates, the refined moment returns, within error, to its zero-field value.
Unlike in CoCl2H2O, the emergence of a field-driven ferromagnetic phase is not preceded by a collective reorientation of only one-third of the chains. The difference in behavior may be due to the distance between anti-aligned chains in the zero-field structures (5.6 Å for CoCl2H2O and 6.2 Å for Li2CoCl4), producing a weaker next-nearest-neighbor (antiferromagnetic) exchange interaction in Li2CoCl4, and thus a lower barrier to overcoming the interaction with applied field. DFT results from the previous work also indicated that for a 3.5 K experimental cell, the moment arrangement was higher in energy than the moment arrangement by only 0.62 meV/atom [18]. Trivially, for an isolated 2.10 moment along paired with an anti-aligned 1.6 T field, the energy gain for flipping the moment is 0.39 meV (). For 5.5 T, the energy gain would be 1.34 meV. The proposed magnetic structure changes are, therefore, energetically reasonable. At 5.5 T, almost no antiferromagnetic reflection intensity remains.
IV.2 Paramagnetic free-ion model
Spin exchange interactions clearly influence the properties of Li2CoCl4 at the magnetic fields probed with neutron diffraction. The IR data, though, includes information not only in those field regions, but also at fields significantly above the moment saturation field, where the Zeeman energy dominates antiferromagnetic spin-spin interactions and the magnetic anisotropy energy [30]. We can, therefore, approximate the spectroscopic behavior of Li2CoCl4 with a paramagnetic spin model at high fields. The paramagnetic resonance energy equals with a -factor range of [, ] depending on the crystallite orientation relative to the magnetic field. We estimated the -factors by fitting high-field data slices with line shape simulations of a spin- system. For the fits (e.g. Fig. 7b), we used the linewidth of the strongest zero-field peak (Fig. 7c). Introducing additional in-plane anisotropy overparameterized the model, with slight variations in the value ( = 0.1) having no significant impact on subsequent analysis.

### v3

@@ Abstract
Insulators containing chains of magnetic transition metal cations provide platforms for probing spin- dynamics and quantum critical behavior. Li2CoCl4 contains edge-sharing CoCl6 octahedra that form chains along the crystallographic axis and orders antiferromagnetically at zero field, but questions remain about its applied-field magnetic structure and the Co2+ spin state. Here, we show with neutron diffraction on a polycrystalline sample how the anti-aligned chains of cobalt moments begin to transition to a ferromagnetic state above 1.6 T. Further, using magnetic resonance absorption measurements and noninteracting spin models, we reveal the strongly anisotropic nature of the Co2+ ion’s -like magnetic behavior ( and ) and its ground state. We, therefore, supply the magnetic structures and anisotropic description needed to explore the dynamics of the field-driven magnetic phases, laying the foundation for further experimental and theoretical studies.
@@ I Introduction
At zero field, Li2CoCl4 orders antiferromagnetically with magnetic space group symmetry. The Co2+ magnetic moments align within edge-sharing CoCl6 chains and anti-align between nearest neighbor chains. The magnetic susceptibility and heat capacity shift with increasing field, leading to three proposed magnetic phase regions separated by transitions near 1.6 T and 3.5 T at 2 K [18]. The magnetic structures of the applied-field regions are unknown. The low-field region is antiferromagnetic, likely similar to the zero-field magnetic ordering. The intermediate-field region may be ferrimagnetic, with chain rotations reminiscent of the structurally similar compound CoCl2H2O [23, 24], though the data were also consistent with a spin-flop reorientation. Finally, in the high-field region, not sharply delineated from the intermediate-field one, the Li2CoCl4 moment chains are likely aligned, but with no well-defined Curie temperature in the magnetic susceptibility.
Here, we use polycrystalline Li2CoCl4 to probe the applied-field regions’ magnetic structures with neutron diffraction, particularly that of the previously ambiguous intermediate-field region. Additionally, we extract several parameters from applied-field infrared spectroscopy data, such as the spectroscopic -factors, to discuss the role of spin-orbit coupling and octahedral distortion on the magnetic properties of Li2CoCl4.
@@ II Materials and methods
We collected neutron diffraction data on the HB-2A beamline at the High Flux Isotope Reactor at Oak Ridge National Laboratory (ORNL). To avoid grain reorientation under applied field, we pressed pellets of Li2CoCl4 for the diffraction experiment. Under argon, we pressed 1.5 g of powder into around 30 1/8” diameter pellets. We then sealed the pellets under vacuum, transported them to ORNL, and loaded them under helium into a 6 mm diameter vanadium canister. The top of the canister was packed with aluminum foil before being hermetically sealed.
@@
We performed magneto-infrared measurements at the infrared (IR) spectroscopy facility at the National High Magnetic Field Laboratory with a 17.5 T vertical-bore superconducting magnet coupled to a Bruker Vertex 80v FTIR spectrometer. In an argon glovebox, a 3–5 mg polycrystalline sample was bonded with n-eicosane and loaded between two polypropylene layers to protect it from oxygen and moisture. We placed the sample in a Voigt geometry so that the incoming IR light was perpendicular to the magnetic field. A composite Si bolometer (IR Labs) collected the transmitted IR light with a spectral range of 10–1100 cm-1 (0.3–33 THz) and a resolution of 0.3 cm-1 (9 GHz). Additionally, we placed a low-pass THz filter (QMC Instruments) with a 120 cm-1 cutoff frequency in front of the sample to increase sensitivity in the far-IR (FIR) range and to minimize radiative heating. IR data were collected near 5 K.
The FIR transmission spectra were divided by a reference spectrum constructed from the maximum transmission at each wavenumber across all spectra measured at different magnetic fields. Such normalized transmission spectra, , are only sensitive to intensity changes induced by the magnetic field and eliminate contributions from magnetic-field-independent vibrational absorption and instrument artifacts. The magnetic resonance absorption spectrum is then calculated as , and results were analyzed using the EPR analysis package EasySpin [28].
III Results and Discussion
@@ III.1 Zero- and low-field magnetic structures
We first confirmed that the zero-field magnetic structure matches that determined previously. The data refines well to the antiferromagnetic structure with (BNS no. 55.363) magnetic space group symmetry (see supplemental material [29]). For the magnetic space group , only reflections where are allowed, whereas the nuclear space group only allows reflections where , making it simple to identify the magnetic ordering reflections. The magnetic structure is commensurate, with the propagation vector k = (1,0,0). The magnetic moments are aligned within the Co2+ chains and anti-aligned between nearest neighboring chains along the nuclear cell’s -plane diagonal. The antiferromagnetic interchain interactions lead to the previously reported cusp in the magnetic susceptibility and the negative Curie-Weiss temperature [18]. The refined moment is 2.10(3) . Though the moment in this study is slightly lower than the previously refined value 2.19(4) , the refinement standard errors are likely underestimated. Since we were able to produce a closer fit to peak intensities in this work, especially nuclear intensity, we will use 2.10 here. The 20 K, zero-field data in the paramagnetic region is included in the supplemental material for comparison [29].
@@
Prior magnetic susceptibility and heat capacity data indicate that the low-field magnetic structure ( 1.6 T) matches the zero-field structure. At 0.8 T, we observe the same magnetic reflections as in the zero-field case, associated with a loss of -centering, but with slightly decreased intensity. Refinements give an unconstrained magnetic moment of 2.02(3) (Fig. 1), indicating some magnetic intensity may become paramagnetic background, though restricting the moment to 2.10 gives a nearly identical fit. With the field increased to 1.5 T, just below the 1.6 T transition to the intermediate-field phase, we can again refine the data well to the structure with a constrained 2.10 moment or a refined moment of 1.96(3) . See the supplemental material [29] for the 1.5 T refinement.
@@ III.2 Field-induced moment rotation
Between the 1.5 T and 2.5 T scans, the antiferromagnetic structure reflections, e.g. (010), (100), and (120), lose significant intensity while the reflections allowed by the -centered nuclear cell, e.g. (110) and (020), gain intensity. The intensity shifts indicate a spin reorientation to a k = 0 magnetic structure. Because of the nuclear cell’s centering, only ferromagnetic arrangements are consistent with k = 0 ordering. Unlike for CoCl2H2O [23, 24], we do not observe new peaks associated with a ferrimagnetic phase where one-third of the chains are anti-aligned. The difference in behavior may be due to the distance between anti-aligned chains in the zero-field structures (5.6 Å for CoCl2H2O and 6.2 Å for Li2CoCl4), producing a weaker next-nearest-neighbor (antiferromagnetic) exchange interaction in Li2CoCl4, and thus a lower barrier to overcoming the interaction with applied field. Instead, in Li2CoCl4 a spin reorientation begins to align the chains of Co2+ magnetic moments. The exact nature of the moment reorientation cannot be determined. If the reorientation involves a spin-flop, as previously proposed, a magnetic field applied along the axis would lead to a temporary “flop” to a hard axis perpendicular to before a strong enough field produced a ferromagnetic arrangement of moments along . Since the magnetic field is not applied along a specific crystallographic direction, we cannot confirm nor repudiate the presence of a spin-flop phase.
The randomly oriented structural grains lead to a slow evolution of the magnetic diffraction intensity with increasing field and a mixture of magnetic phases. For some of the randomly oriented structural grains in the sample, the magnetic field will align with the crystallographic axis preferred by the magnetic anisotropy, leading to the lowest-energy ferromagnetic arrangement and coherent magnetic scattering intensity. In contrast, for grains where the magnetic field is not aligned with the anisotropy-preferred direction, the competing perturbations will lead to moment canting off the anisotropy-preferred axis toward the field, explaining the previously observed net magnetization for a polycrystalline sample [18]. The magnetic scattering reflects the random distribution of these grains with canted moments and thus their randomly distributed moment orientations.
Several ferromagnetic models can successfully capture the additional magnetic intensity at high fields. The simplest model uses the maximal magnetic space group (BNS no. 65.485), which constrains the cobalt magnetic moments along the crystallographic axis. Additional maximal magnetic groups are considered in the supplemental material [29]. Fig. 2 shows the and magnetic unit cells along with the nuclear cell. Considering lower-symmetry, commensurate magnetic structures with canted moments does not improve fits to the ferromagnetic intensity. For example, the 2 space group symmetry (BNS no. 12.62) is the highest symmetry k = 0 magnetic space group possible for Li2CoCl4 that allows for moment components along and . Refinements to the 5.5 T data using 2 ( = 1.5(2) , = 1.7(1) , ) gave a similar fit to the simpler model [29]. Likewise, an arrangement with magnetic symmetry (BNS no. 2.4) allows for moment components along all three crystallographic directions and gives a similar fit (). While there is ambiguity in the refinements regarding the , , and magnetic symmetries, the scenario is likely more consistent with the polycrystalline sample’s evolution due to the previously discussed randomness of the structural grains’ orientations relative to the applied magnetic field, leading to an average magnetic structure with moment components along all three crystallographic directions.
The best refinement models above 1.6 T involve both the antiferromagnetic magnetic phase and a ferromagnetic k = 0 structure. Using the simplest model ( symmetry) allows us to track the evolution of the magnetic structure of Li2CoCl4 at magnetic fields greater than 1.6 T by refining the data with contributions from the antiferromagnetic structure and the ferromagnetic structure. The moment is constrained to be equal in the two magnetic phases when both are refined. The full range refinements for the 2.5, 3.5, 4.5, and 5.5 T data are shown in Fig. S4 [29]. Table 1 contains the refined phase fractions of the and magnetic structures and the refined cobalt magnetic moment for each magnetic field.
To highlight the changing magnetic contributions with increasing field, Fig. 3 compares the 0.8, 2.5, and 5.5 T refinements, corresponding to the previously described low-, intermediate-, and high-field regions. Since the intermediate-field region involves the greatest competition between the antiferromagnetic and ferromagnetic interactions, the refined magnetic moment is significantly reduced for the 2.5 and 3.5 T data. Likewise, while most of the refinements are qualitatively identical if the cobalt magnetic moment is constrained to the zero-field value of 2.10 , the 2.5 and 3.5 T refinements significantly change (Fig. S5 [29]). At higher fields where the ferromagnetic phase dominates, the refined moment returns, within error, to its zero-field value, and constraining the moment to 2.10 does not influence the fit.
@@
The emergence of the ferromagnetic phase with increasing field can be monitored more closely by tracking the relatively high intensity (010) reflection associated with the antiferromagnetic phase (Fig. 4). Above 1.6 T, the (010) intensity drops until the intensity levels off near 4.5 T. The drop in (010) reflection intensity at 1.6 T matches the slope change of the magnetization curve, the susceptibility’s leveling off below the critical temperature, and the emergence of a hump in the magnetic heat capacity alongside a sharper peak [18]. The suppression of the heat capacity peak along with the emergence of a hump, therefore, is due to the reduction of antiferromagnetic regions and the short-range aligning of chains in newly formed ferromagnetic regions of the randomly oriented polycrystalline sample. While tracking the (010) reflection, we also captured changes in lower intensity reflections with the other diffractometer detectors: the antiferromagnetic (100) reflection and the nuclear/ferromagnetic (131) reflection. Both show the same trend of increasing ferromagnetic intensity and decreasing antiferromagnetic intensity above 1.5–1.6 T [29].
@@
By the time the field reaches 5.5 T, the (010) reflection intensity has approached the experimental background, the majority of the sample is no longer antiferromagnetic, and the magnetic anisotropy likely favors an arrangement of moments along the axis. Using the previous DFT results for a 3.5 K experimental cell, the moment arrangement is 0.62 meV/atom higher in energy than the moment arrangement [18], so, trivially, for an isolated 2.10 moment along paired with an anti-aligned 5.5 T field, the energy gain for flipping the moment is 1.34 meV (). The proposed magnetic structure change is, therefore, energetically reasonable. In the ferromagnetic-dominated field region, we tracked the (110) transition with increasing temperature (Fig. S8). Up to 20 K, we do not observe a sharp drop or leveling off of the intensity but rather a nearly linear decrease in intensity. Similarly, the magnetic susceptibility at high field, even up to 7 T (Fig. S8), does not show the typical sharp rise of a paramagnetic to ferromagnetic transition, suggesting that the high-field magnetic ordering is driven by field overcoming the material’s interchain antiferromagnetic exchange interactions instead of by dominant interchain ferromagnetic exchange interactions.
III.3 Spin model from high-field IR spectroscopy
To explore the spin dynamics of Li2CoCl4 under high applied fields, we collected IR spectroscopy data up to 17.5 T. At lower energies within the detection range, the normalized magnetic resonance absorption contains a broad continuum (Fig. 5a). The broad continuum reveals the strongly anisotropic nature of the magnetic moment of the Co2+ cation in the Li2CoCl4 powder sample. At magnetic fields significantly above the moment saturation field, the Zeeman energy dominates the antiferromagnetic spin-spin interactions, and spin dynamics converge toward those of free-ion spin precession [30]. For many low-dimensional cobalt compounds with octahedral ligand environments, the magnetic properties are well-described by an effective spin- model due to a well-separated ground state Kramers doublet [31, 32, 33, 34]. Therefore, as a first approximation, we use a noninteracting spin- framework to extract -factors from the high-field IR data.
III.3.1 Noninteracting spin- model
In the free spin- framework, an applied magnetic field splits the and states (Zeeman splitting) by the energy , where can be directly observed as a peak in the absorption spectrum. The parameter is the effective -factor, whose value depends on the orientation of the magnetic field relative to the anisotropy axis. Due to the randomly oriented grains in the polycrystalline sample, the absorption associated with the Zeeman splitting is distributed over a frequency range that broadens with magnetic field, resulting in the continuum in Fig. 5a. The asymmetric lineshape of the absorption spectrum (Fig. 5b) indicates a system with axial symmetry, and indeed the CoCl6 octahedra in Li2CoCl4 are slightly compressed along the axis [ratio of the Co-Cl bond length along to the length within the plane of 0.986(2)], producing a tetragonal distortion of the octahedra. The continuum is bounded by resonance conditions and with more intensity on the side [35], where and are relative to the compressed CoCl6 axis. Since the continuum edges at the resonance energies associated with and are smeared by peak broadening, we simulated the absorption spectrum using a powder-averaged lineshape and varied the -factors to obtain a match (Fig. 5b). The fit anisotropic -factors are = 2.77 and = 5.23. The corresponding resonance lines confine the magnetic absorption spectrum for a free cobalt ion, but the experimental data diverge from the lines below 10 T, where exchange interactions produce a more complex picture of collective spin excitations (Fig. 5a).